OpenStudy (anonymous):

How much interest will be paid in 1 year for a loan of $2500 at 9 1/2% simple interest?

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

can you identify the principal and the interest rate?

OpenStudy (texaschic101):

Interest = Principle * rate * time ..... I = PRT P = 2500 R = 9 1/2 % = 9.5% = 0.095 T = 1 Now we sub into the formula I = (2500)(0.095)(1) now just multiply and what do you get ?

OpenStudy (anonymous):

237.5

OpenStudy (anonymous):

that's correct...add that number to the original number to get the total .:

OpenStudy (anonymous):

237.5+2500=2,737.5